How many moles of Z (aq) must be added to an empty 300 mL container to achieve

equilibrium molarity for R (aq) of 0.3 mol/L given the following system

2Q (aq) + R (aq) ⇌ 3Z (aq),

Where the K value is 3.2 x 10-6?


1
Expert's answer
2021-06-10T03:33:46-0400

0.3 = 3.2 × 10-6 × M/0.3

M × 3.2 × 10-6 = 0.3 × 0.3

M = 0.09/3.2 × 10-6

= 2.8125 moles